y=2x=1;x=-1,x=-2,y=3 find the solutions of the given equation using the given values of x and y. i am not quite sure how to do this could you help me please?
Answer by stanbon(75887)

y=2x+1
When x=-1, y = 2(-1)+1 = -2+1 = -1
---------------------------------------
When x=-2, y = 2(-2)+1 = -4+1 = -3
---------------------------------------
When y=3, 3 = 2x+1; 2x=2, x = 1
=========
Cheers,
Stan H.